# brick-splitting Torsion Pairs and Trim Lattices

### v1

@@ 3.1. Basic properties
$0\to X\stackrel{{\scriptstyle u}}{{\to}}E\stackrel{{\scriptstyle v}}{{\to}}Y\to 0$
@@
Let $X\in\mathcal{B}$ or $X\in\mathcal{C}$. We show that $X$ is rigid. By assumption, there is a left finite or right finite semibrick $\mathcal{S}$ such that $X\in\mathcal{S}$. By [As1, Theorem 2.27] and its dual, we have that $\operatorname*{Filt}\mathcal{S}$ is equivalent to the module category of some finite dimensional algebra $\Gamma$. We can take an idempotent $e\in\Gamma$ such that $\operatorname*{Filt}X\simeq\operatorname*{mod}\Gamma/\langle e\rangle$, since $\operatorname*{Filt}X$ is a Serre subcategory of $\operatorname*{Filt}\mathcal{S}$. Denote by $Y$ the indecomposable projective object in $\operatorname*{Filt}X$. Then, $\operatorname{Ext}^{1}_{A}(Y,X)=0$. Since $X$ is a subobject of $Y$, we have a monomorphism $\iota:X\to Y$. Because $A$ is hereditary, this gives rise to an epimorphism $0=\operatorname{Ext}^{1}_{A}(Y,X)\to\operatorname{Ext}^{1}_{A}(X,X)$, showing that $X$ is rigid.
@@
We note that, unlike $\mathcal{T}$ and $\mathcal{T}^{\prime}$, a simple computation shows that the torsion class labeled by $\{{\begin{smallmatrix}2\\ 2\end{smallmatrix}}\}$ is not brick-splitting. More precisely, if $\mathcal{T}^{\prime\prime}:=\operatorname{Fac}({\begin{smallmatrix}2\\ 2\end{smallmatrix}})$, and ′′$\mathcal{F}^{\prime\prime}:=\mathcal{T}^{\prime\prime\perp}$, then the brick ${\begin{smallmatrix}1\\ 2\end{smallmatrix}}$ does not belong to $\mathcal{T}^{\prime\prime}$, nor $\mathcal{F}^{\prime\prime}$.
@@
$\operatorname*{mod}A$$\{{\begin{smallmatrix}&1&\\ 2&&1\end{smallmatrix}}\}$$\{{\begin{smallmatrix}1\\ 1\end{smallmatrix}}\}$00$\{{\begin{smallmatrix}2\\ 2\end{smallmatrix}}\}$
@@
Using Proposition 3.7, we obtain the following interesting result. As previously defined in Section 1, $A$ is called a fully cyclic algebra if for any idempotent $e$ of $A$, from $eA(1-e)=0$ we get $e=0$ or $e=1$. Now we prove Corollary 1.3.
@@
Assume that there is a non-trivial brick-splitting torsion pair $(\mathcal{T},\mathcal{F})$ in $\operatorname*{mod}A$. Then, each of $\operatorname*{sim}A\cap\mathcal{F}$ and $\operatorname*{sim}A\cap\mathcal{T}$ is non-empty. If $\Omega_{0}$ denotes the vertex set of $\Omega$, then it follows from Proposition 3.7 that we can partition $\Omega_{0}$ into two non-empty subsets $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ such that there is no arrow from a vertex of $\mathcal{S}_{1}$ to a vertex of $\mathcal{S}_{2}$. If $e$ denotes the idempotent corresponding to the vertices belonging to $\mathcal{S}_{1}$, we have $(1-e)Ae=0$. Since $e\neq 0$ and $e\neq 1$, then $A$ is not fully cyclic. This is the desired contradiction, and we conclude the first assertion.
Now, we check that any connected self-injective algebra is fully cyclic. If we assume otherwise, then we get $eA(1-e)=0$ for some idempotent $0\neq e\neq 1$. Hence, the projective module $A(1-e)$, which has to be injective, is supported only on the vertices in $\Omega_{0}$ corresponding to $1-e$. This yields that $D((1-e)A)\cong A(1-e)$. By the connectedness of $A$, we get $(1-e)A(1-e)=A$, a contradiction. ∎
@@
3.2. Characterizations
Proposition 3.12
  1. (1)
    $(\mathcal{T},\mathcal{T}^{\perp})$ is a brick-splitting torsion pair;
  2. (2)
Proof.
Assume that (1) holds. Consider $X\in\operatorname*{brick}A$. Since $(\mathcal{T},\mathcal{T}^{\perp})$ is brick-splitting, we get $X\in\mathcal{T}$ or $X\in\mathcal{T}^{\perp}$. In the first case, $\mathcal{T}(X)$ is a torsion class in the interval $[0,\mathcal{T}]$, and we have $X$ is the labeling brick of the minimal inclusion $\mathcal{T}(X)\cap^{\perp}X\subset\mathcal{T}(X)$. In the second case, if $X\in\mathcal{T}^{\perp}$, then $X$ is the labeling brick of the minimal inclusion $\mathcal{F}(X)\cap X^{\perp}\subset\mathcal{F}(X)$ of torsion-free classes, all included in $\mathcal{T}^{\perp}$. This translates to $X$ being a brick-labeling of the corresponding inclusion ${}^{\perp}X\subset{{}^{\perp}(\mathcal{F}(X)\cap X^{\perp}})$ of torsion classes in $[\mathcal{T},\operatorname*{mod}A]$.
Now, assume that $(2)$ holds. Let $X\in\operatorname*{brick}A$. If $X$ is a labeling brick between two torsion classes in $[0,\mathcal{T}]$, then obviously $X\in\mathcal{T}$. Suppose that $X$ is the labeling brick between some two adjacent torsion classes $\mathcal{T}_{1}\subset\mathcal{T}_{2}$ in the interval $[\mathcal{T},\operatorname*{mod}A]$. This implies that $X\in\mathcal{T}_{1}^{\perp}\subseteq\mathcal{T}^{\perp}$. ∎
Remark 3.13
For brick-finite algebras, the above proposition gives an easy criterion to verify whether a given torsion class admits a brick-splitting torsion pair. For instance, in Example 3.5, one can check that every $X\in\operatorname*{brick}A$ appears on the left chain of the lattice $\operatorname*{tors}A$ (see Figure 1). Consequently, every torsion class on this chain gives rise to a brick-splitting torsion pair. In contrast, the labeling bricks on the other maximal chain on the right side consists of only two bricks, whereas $\operatorname*{brick}A$ has three elements. Thus, we can immediately conclude that $\mathcal{T}^{\prime\prime}=\operatorname{Fac}(\begin{smallmatrix}2\\ 2\end{smallmatrix})$ is not a brick-splitting torsion class. In particular, ${\begin{smallmatrix}1\\ 2\end{smallmatrix}}\in\operatorname*{brick}A$ does not appear in $[0,\mathcal{T}^{\prime\prime}]$ nor $[\mathcal{T}^{\prime\prime},\operatorname*{mod}A]$.
Corollary 3.14
Proof.
Proposition 3.15
  1. (1)
    $(\mathcal{T},\mathcal{T}^{\perp})$ is a brick-splitting torsion pair;
  2. (2)
Proof.
$0\to T\to L\to F\to 0$
for $L$ where $T\in\mathcal{T}$ and $F\in\mathcal{T}^{\perp}$. Since $L$ is a minimal extending brick for $\mathcal{Y}$, we know that every proper quotient of $L$ has to be in $\mathcal{Y}$. So if $T$ is non-zero (we know that $T\neq L$), then $F\in\mathcal{Y}$, making $L\in\mathcal{Y}\vee\mathcal{T}$, a contradiction. Therefore, $T=0$ and $L\in\mathcal{T}^{\perp}$.
For the other direction, assume that $\mathcal{T}$ is brick-splitting. Observe that it follows from [DI+, Theorem 1.3] that $\operatorname*{tors}A$ is weakly atomic. It follows from [TW] that in order to prove that $\mathcal{T}$ is left modular, it suffices to prove that for two torsion classes $\mathcal{V}\subsetneq\mathcal{U}$ forming a cover relation, we have
Observe that the containment from left to right is clear. Now, let $L$ be the labeling brick corresponding to the cover relation $\mathcal{V}\subsetneq\mathcal{U}$. Since $\mathcal{T}$ is brick-splitting, either $L\in\mathcal{T}$ or $L\in\mathcal{T}^{\perp}$. If $L\in\mathcal{T}$, we see that
$\mathcal{V}\vee\mathcal{T}=\mathcal{V}\vee(\mathcal{T}(L)\vee\mathcal{T})=(\mathcal{V}\vee\mathcal{T}(L))\vee\mathcal{T}=\mathcal{U}\vee\mathcal{T}.$
In this case, the right-hand-side of ()(*)( ∗ ) becomes $\mathcal{U}$. Since the left-hand-side contains both $L$ and $\mathcal{V}$, then it has to contain $\mathcal{U}$. This proves that ()(*)( ∗ ) holds. Assume now that $L\in\mathcal{T}^{\perp}$. Observe that $\mathcal{V}=\mathcal{U}\cap{{}^{\perp}L}$. Hence,
$\mathcal{V}\cap\mathcal{T}=(\mathcal{U}\cap{{}^{\perp}L})\cap\mathcal{T}=\mathcal{U}\cap(^{\perp}L\cap\mathcal{T})=\mathcal{U}\cap\mathcal{T}.$
In this case, the left-hand-side of ()(*)( ∗ ) is $\mathcal{V}$. Now, observe that both $\mathcal{T}$ and $\mathcal{V}$ are contained in ${}^{\perp}L$. Hence,
$\mathcal{U}\cap(\mathcal{T}\vee\mathcal{V})\subseteq\mathcal{U}\cap{{}^{\perp}L}=\mathcal{V}.$
Therefore, ()(*)( ∗ ) also holds in this case. ∎
@@ 4.1. Characterizations of brick-directedness
As introduced in Section 1, in this work we generalize the above classical notion of representation-directedness. In particular, a cycle $X_{0}\xrightarrow{f_{1}}X_{1}\xrightarrow{f_{2}}\cdots\xrightarrow{f_{m-1}}X_{m-1}\xrightarrow{f_{m}}X_{m}=X_{0}$ in $\operatorname*{mod}A$ is called a brick-cycle if $X_{i}\in\operatorname*{brick}A$, for all $0\leq i\leq m$. We say an algebra $A$ is brick-directed if there exists no brick-cycle in $\operatorname*{mod}A$. From the definition, it is immediate that each representation-directed algebra is brick-directed. However, there are many brick-directed algebras that are not representation-directed.
@@
As recalled in Theorem 2.7, from [KD] we know that each chain of bricks of $A$ gives rise to a chain of torsion classes in $\operatorname*{mod}A$, and this correspondence induces a bijection between the equivalence classes of maximal chains of bricks and the set of maximal chains of torsion classes in $\operatorname*{tors}A$. In fact, an interesting characterization of brick-directed algebras can be given in terms of their maximal chain of torsion classes and the notion of brick-splitting torsion pairs.
Theorem 4.7
For an algebra $A$, the following are equivalent:
  1. (1)
    $A$ is brick-directed;
  2. (2)
    There exists a total order $J$ on $\operatorname*{brick}A=\{B_{j}\}_{j\in J}$ with $\operatorname{Hom}_{A}(B_{i},B_{j})=0$, whenever $i<j$ in $J$.
  3. (3)
Proof.
Assume that (1) holds. Consider the transitive closure $\leq$ of the relation $\sim$ in $\operatorname*{brick}A$ given by $X\sim Y$ if $\operatorname{Hom}_{A}(X,Y)\neq 0$. By the brick-directedness, the relation $\leq$ is anti-symmetric, hence is an order relation. Now, it is possible to well-order $\operatorname*{brick}A$ by $\preceq$ so that $X\leq Y$ implies $X\preceq Y$. The opposite order to $\preceq$ is the order that gives us (2).
Assume that (2) holds. We consider the set $\mathcal{I}$ of all down-closed subsets of $J$. Then $\mathcal{I}$ can be totally ordered by inclusion. For $I\in\mathcal{I}$, let $\mathcal{T}_{I}$ be the smallest torsion class that contains all $B_{i}$ for $i\in I$. If $j\in J$ is such that $i<j$ for all $i\in I$, then it is clear that $\operatorname{Hom}_{A}(\mathcal{T}_{I},B_{j})=0$ and hence that $B_{j}\in\mathcal{T}_{I}^{\perp}$. This proves that each $\mathcal{T}_{I}$ is a torsion class that is brick-splitting. We consider the chain $\{\mathcal{T}_{I}\}_{I\in\mathcal{I}}$. Assume that this chain is not maximal and let $\mathcal{T}$ be a torsion class that can be added to this chain. Let $K\subseteq J$ corresponding to the bricks in $\mathcal{T}$. Then $K$ does not form a down closed set. Hence, there exist $B_{i},B_{j}\in\operatorname*{brick}A$ such that $B_{j}\in\mathcal{T}$ while $B_{i}\not\in\mathcal{T}$, with $i<j$. Now, the torsion class $\mathcal{T}_{\leq i}$ has to be comparable to $\mathcal{T}$. Since $B_{i}\not\in\mathcal{T}$, we have $\mathcal{T}\subseteq\mathcal{T}_{\leq i}$, which leads to the contradiction that $B_{j}\in\mathcal{T}_{\leq i}$. Hence $\{\mathcal{T}_{I}\}_{I\in\mathcal{I}}$ is a maximal chain of brick-splitting torsion classes. This proves (3).
Finally, assume that (3) holds. Suppose to the contrary that (1) does not hold, and let $B_{1}\to B_{2}\to\cdots\to B_{n}\to B_{1}$ be a brick-cycle, where $n\geq 2$. Consider $\mathcal{T}$ the union of all torsion classes in the chain that contain no $B_{i}$. Clearly $\mathcal{T}$ is again a torsion class that is brick-splitting, and is comparable to the torsion classes of the chain, hence has itself to be in the chain. Similarly, we let $\mathcal{T}^{\prime}$ be the intersection of all torsion classes in the chain that contain all $B_{i}$. Then $\mathcal{T}^{\prime}$ is a torsion class that is brick-splitting, and is comparable to the torsion class of the chain, hence has to be in the chain. Now, we consider the inclusion $\mathcal{T}\subsetneq\mathcal{T}^{\prime}$. If this is not a cover relation, there is a torsion class $\mathcal{T}^{\prime\prime}$ in the chain that contains a non-empty proper subset of $\{B_{1},\ldots,B_{n}\}$. In particular, there is $i$ such that $B_{i}\in\mathcal{T}^{\prime\prime}$ while $B_{i+1}\not\in\mathcal{T}^{\prime\prime}$, where indices are taken modulo $n$. By the brick-splitting property, $B_{i+1}\in(\mathcal{T}^{\prime\prime})^{\perp}$ so that $\operatorname{Hom}_{A}(B_{i},B_{i+1})=0$, a contradiction. Hence, the inclusion $\mathcal{T}\subsetneq\mathcal{T}^{\prime}$ is a cover relation. In particular, it follows from [DI+, Theorem 3.3(b)] that there is a unique brick in $\mathcal{T}^{\prime}\cap\mathcal{T}^{\perp}$. Since $\mathcal{T}$ is brick-splitting, we have $\{B_{1},\ldots,B_{n}\}\subseteq\mathcal{T}^{\prime}\cap\mathcal{T}^{\perp}$ so that $n=1$, a contradiction. ∎
Corollary 4.8
If $A$ is brick-directed, then $A$ is weakly triangular.
Proof.
Remark 4.9
We remark that weakly triangular algebras appeared in the study and classification of geometrically irreducible algebras (over algebraically closed fields). These are the algebras for which every module variety $\operatorname*{mod}(A,\underline{d})$ is irreducible, where the dimension vector $\underline{d}\in\mathbb{Z}^{n}_{\geq 0}$ varies. In fact, it is shown that every geometrically irreducible algebras is weakly triangular. For more details and new results, see [BZ] and the references therein.
We now give a more combinatorial realization of brick-directed algebras. First, observe that, to any algebra, one can explicitly associate a quiver that captures the interactions between the bricks and hence the brick-directedness of the algebra under consideration. In particular, for a connected algebra $A$, by $Q^{b}(A)$ we denote the brick-quiver of $A$, constructed as follows: The vertices of $Q^{b}(A)$ are in bijection with elements of $\operatorname*{brick}A$, and for each pair of distinct (hence non-isomorphic) modules $X$ and $Y$ in $\operatorname*{brick}A$, we put an arrow from $X$ to $Y$ if $\operatorname{Hom}_{A}(X,Y)\neq 0$. For an explicit example, see Example 4.14.
Proposition 4.10
  1. (1)
    $Q^{b}(A)$ is connected and has no loops.
  2. (2)
  3. (3)
  4. (4)
  5. (5)
Proof.
  1. (1)
    $0\to S_{j}\to M\to S_{i}\to 0$
  2. (2)
  3. (3)
  4. (4)
    This directly follows from the definition.
  5. (5)
As an immediate consequence of the preceding proposition, we obtain the following characterization of brick-directed algebras in terms of their brick-quivers.
Corollary 4.11
Let $A$ be a connected algebra. Then, $A$ is brick-directed if and only if the brick-quiver $Q^{b}(A)$ is acyclic.
Remark 4.12
As observed in [MP2], from some earlier results of [As1], it follows that for any brick-finite algebra $A$ of rank $n$, there is no set of pairwise Hom-orthogonal modules of size $n+1$. In particular, for such an algebra, the maximal size of an independent set in $Q^{b}(A)$ is $n$, that is, the independence number of $Q^{b}(A)$ is $n$ and is attained exactly for the set of vertices in $Q^{b}(A)$ corresponding to the simple $A$-modules (see [MP2, Theorem A]). In light of the Semibrick conjecture (see [MP2, Conjecture 2.2]), it is expected that the converse of the latter statement is true, that is, the independence number of $Q^{b}(A)$ is $n$ if and only if $A$ is brick-finite. This is already known to be the case for various families of algebras, but it remains open in full generality (for more details, see [MP2] and references therein).

### v2

@@ 3.3. Basic properties
$0\to X\stackrel{{\scriptstyle u}}{{\to}}E\stackrel{{\scriptstyle v}}{{\to}}Y\to 0.$
@@
We show that each $X\in\mathcal{B}\cup\mathcal{C}$ is rigid. Let $\mathcal{S}:=\mathcal{B}$ if $X\in\mathcal{B}$, and $\mathcal{S}:=\mathcal{C}$ if $X\in\mathcal{C}$. By [As1, Theorem 2.27] and its dual, we have that $\operatorname*{Filt}\mathcal{S}$ is equivalent to the module category of some finite dimensional algebra $\Gamma$. We can take an idempotent $e\in\Gamma$ such that $\operatorname*{Filt}X\simeq\operatorname*{mod}\Gamma/\langle e\rangle$, since $\operatorname*{Filt}X$ is a Serre subcategory of $\operatorname*{Filt}\mathcal{S}$. Denote by $Y$ the indecomposable projective object in $\operatorname*{Filt}X$. Then, $\operatorname{Ext}^{1}_{A}(Y,X)=0$. Since $Y\in\operatorname*{Filt}X$, we have a monomorphism $\iota:X\to Y$. Because $A$ is hereditary, this gives rise to an epimorphism $0=\operatorname{Ext}^{1}_{A}(Y,X)\to\operatorname{Ext}^{1}_{A}(X,X)$, showing that $X$ is rigid.
@@
We note that, unlike $\mathcal{T}$ and $\mathcal{T}^{\prime}$, a simple computation shows that the torsion class labeled by $\{{\begin{smallmatrix}2\\ 2\end{smallmatrix}}\}$ is not brick-splitting. More precisely, if $\mathcal{T}^{\prime\prime}:=\operatorname{Fac}({\begin{smallmatrix}2\\ 2\end{smallmatrix}})$, and $\mathcal{F}^{\prime\prime}:=\mathcal{T}^{\prime\prime\perp}$, then the brick $\begin{smallmatrix}1\\ 2\end{smallmatrix}$} does not belong to $\mathcal{T}^{\prime\prime}$, nor $\mathcal{F}^{\prime\prime}$.
@@
[…]
@@
Using Proposition 3.14, we obtain the following interesting result. As previously defined in Section 1, $A$ is called a fully cyclic algebra if for any idempotent $e$ of $A$, from eA(1e)=0eA(1-e)=0 we get $e=0$ or $e=1$. Now we prove Corollary 1.3.
@@
Assume that there is a non-trivial brick-splitting torsion pair $(\mathcal{T},\mathcal{F})$ in $\operatorname*{mod}A$. Then, each of $\operatorname*{sim}A\cap\mathcal{F}$ and $\operatorname*{sim}A\cap\mathcal{T}$ is non-empty. If $\Omega_{0}$ denotes the vertex set of $\Omega$, then it follows from Proposition 3.14 that we can partition $\Omega_{0}$ into two non-empty subsets $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ such that there is no arrow from a vertex of $\mathcal{S}_{1}$ to a vertex of $\mathcal{S}_{2}$. If $e$ denotes the idempotent corresponding to the vertices belonging to $\mathcal{S}_{1}$, we have (1e)Ae=0(1-e)Ae=0. Since $e\neq 0$ and $e\neq 1$, then $A$ is not fully cyclic. This is the desired contradiction, and we conclude the first assertion.
Now, we check that any connected self-injective algebra is fully cyclic. If we assume otherwise, then we get eA(1e)=0eA(1-e)=0 for some idempotent $0\neq e\neq 1$. Hence, the projective module A(1e)A(1-e), which has to be injective, is supported only on the vertices in $\Omega_{0}$ corresponding to 1e1-e. This yields that $D((1-e)A)\cong A(1-e)$. By the connectedness of $A$, we get (1e)A(1e)=A(1-e)A(1-e)=A, a contradiction. ∎
@@ 4.1. Characterizations of brick-directedness
As introduced in Section 1, in this work we generalize the above classical notion of representation-directedness. In particular, a cycle $X_{0}\xrightarrow{f_{1}}X_{1}\xrightarrow{f_{2}}\cdots\xrightarrow{f_{m-1}}X_{m-1}\xrightarrow{f_{m}}X_{m}=X_{0}$ in $\operatorname*{mod}A$ is called a brick-cycle if $X_{i}\in\operatorname*{brick}A$, for all $0\leq i\leq m$. We say an algebra $A$ is brick-directed if there exists no brick-cycle in $\operatorname*{mod}A$. From the definition, it is immediate that each representation-directed algebra is brick-directed. However, there are many brick-directed algebras that are not representation-directed (see Section 4.3).
As recalled in Theorem 2.6, from [KD] we know that each chain of bricks of $A$ gives rise to a chain of torsion classes in $\operatorname*{mod}A$, and this correspondence induces a bijection between the equivalence classes of maximal chains of bricks and the set of maximal chains of torsion classes in $\operatorname*{tors}A$. In fact, an interesting characterization of brick-directed algebras can be given in terms of their maximal chain of torsion classes and the notion of brick-splitting torsion pairs. We recall that $Q^{b}(A)$ denotes the brick-quiver of the algebra $A$, as introduced in Section 3.2.
Theorem 4.1
For an algebra $A$, the following are equivalent:
  1. $(i)$
    AA is brick-directed;
  2. $(ii)$
    $Q^{b}(A)$ is acyclic.
Moreover, in this case, there is a bijection between the following objects.
  1. (1)
    A total order $J$ on $\operatorname*{brick}A=\{B_{j}\}_{j\in J}$ with $\operatorname{Hom}_{A}(B_{i},B_{j})=0$, whenever $i<j$ in $J$.
  2. (2)
    A maximal chain $\{\mathcal{T}_{i}\}_{i\in I}$ in $\operatorname*{tors}A$ such that every $\mathcal{T}_{i}$ is a brick-splitting torsion class.
Proof.
From the construction of the brick quiver $Q^{b}(A)$, it is clear that $(i)$ and $(ii)$ are equivalent. Hence, we construct a map between objects from $(1)$ to objects from $(2)$ as follows. We consider a totally ordered set $J$ where $\operatorname*{brick}A=\{B_{j}\}_{j\in J}$ with $\operatorname{Hom}_{A}(B_{i},B_{j})=0$, whenever $i<j$ in $J$. We consider the set $\mathcal{I}$ of all down-closed subsets of $J$. Then $\mathcal{I}$ can be totally ordered by inclusion. For $I\in\mathcal{I}$, let $\mathcal{T}_{I}$ be the smallest torsion class that contains all $B_{i}$, for $i\in I$. The map is
$(*):\quad\{B_{j}\}_{j\in J}\mapsto\{\mathcal{T}_{I}\}_{I\in\mathcal{I}}.$
If $j\in J$ is such that $i<j$ for all $i\in I$, then it is clear that $\operatorname{Hom}_{A}(\mathcal{T}_{I},B_{j})=0$, hence $B_{j}\in\mathcal{T}_{I}^{\perp}$. This proves that each $\mathcal{T}_{I}$ is a brick-splitting torsion class. Assume that the constructed chain $\{\mathcal{T}_{I}\}_{I\in\mathcal{I}}$ is not maximal and let $\mathcal{T}$ be a torsion class that can be added to properly extend this chain. Let $K\subseteq J$ correspond to the bricks in $\mathcal{T}$. From our assumptions, it follows that $K$ does not form a down closed set. Hence, there exist $B_{i},B_{j}\in\operatorname*{brick}A$ such that $B_{j}\in\mathcal{T}$ while $B_{i}\not\in\mathcal{T}$, with $i<j$. Now, the torsion class $\mathcal{T}_{\leq i}$ has to be comparable to $\mathcal{T}$. Since $B_{i}\not\in\mathcal{T}$, we have $\mathcal{T}\subseteq\mathcal{T}_{\leq i}$, which leads to the contradiction $B_{j}\in\mathcal{T}_{\leq i}$. Hence, $\{\mathcal{T}_{I}\}_{I\in\mathcal{I}}$ is a maximal chain of brick-splitting torsion classes.
To prove that the above assignment is bijective, we construct an inverse. Let $\{\mathcal{T}_{i}\}_{i\in I}$ in $\operatorname*{tors}A$ be a maximal chain such that every $\mathcal{T}_{i}$ is a brick-splitting torsion class. Given a brick $B$, it follows from the maximality of the chain that there is a unique $i_{B}\in I$ with the property that $B\in\mathcal{T}_{i_{B}}$ and for $j<i_{B}$, we have $B\not\in\mathcal{T}_{j}$. Indeed, one takes $i_{B}=\bigwedge_{B\in\mathcal{T}_{i}}i$. If we consider the subchain $\{\mathcal{T}_{i_{B}}\}_{B\in\operatorname*{brick}A}$, this induces a total order $\leq$ on $\operatorname*{brick}A$. If $U,V\in\operatorname*{brick}A$ and $U<V$ with respect to this order, then $\mathcal{T}_{i_{U}}\subset\mathcal{T}_{i_{V}}$ and it follows that $V\not\in\mathcal{T}_{i_{U}}$. Since $\mathcal{T}_{i_{U}}$ is brick splitting, we get $V\in\mathcal{T}_{i_{U}}^{\perp}$ so that $\operatorname{Hom}_{A}(U,V)=0$. Note that with the notation $B_{i_{B}}=B$, the totally ordered set $\{i_{B}\mid B\in\operatorname*{brick}A\}$ gives us a way to label the bricks.
We finish the proof by checking that the assignment
$\{\mathcal{T}_{i}\}_{i\in I}\mapsto\{B_{j}\}_{j\in\{i_{B}\mid B\in\operatorname*{brick}A\}}$
provides an inverse of ()(*). Given $\{B_{j}\}_{j\in J}$, the map ()(*) assigns to it the chain $\{\mathcal{T}_{I}\}_{I\in\mathcal{I}}$ which is such that $i_{B}=\{j\in J\mid B_{j}\leq B\}$. Clearly, for $U,V\in\operatorname*{brick}A$, we have $U\leq V$ if and only if $i_{U}\subseteq i_{V}$, which proves that the second map is a left inverse to ()(*). Conversely, given a maximal chain $\{\mathcal{T}_{i}\}_{i\in I}$ of brick-splitting torsion classes, the second map assigns to it a total order on the bricks belonging to $\{i_{B}\mid B\in\operatorname*{brick}A\}$. Given $i\in I$, if $B$ is a brick in $\mathcal{T}_{i}$, then $\mathcal{T}_{i_{B}}\subseteq\mathcal{T}_{i}$. Because every torsion class is uniquely determined by the bricks it contains, we get $\mathcal{T}_{i}=\bigvee_{B\in\mathcal{T}_{i}}\mathcal{T}_{i_{B}}$, which means that $i=\bigvee_{i_{B}\leq i}i_{B}$. This yields that the second map is also a right inverse to ()(*). ∎
Corollary 4.2
If $A$ is brick-directed, then $A$ is weakly triangular.
Proof.
Remark 4.3
We note that weakly triangular algebras have also appeared in the study and classification of geometrically irreducible algebras (over algebraically closed fields). These are the algebras for which every module variety $\operatorname*{mod}(A,\underline{d})$ is irreducible, where the dimension vector $\underline{d}\in\mathbb{Z}^{n}_{\geq 0}$ varies. In fact, it is shown that every geometrically irreducible algebras is weakly triangular. For more details and new results, see [BZ] and the references therein.
4.2. Basic properties and examples of brick-directed algebras
We start this subsection with an example of a brick-directed algebra that is not representation-directed.